# Monte Carlo evaluations of gamma-ray and radio pulsar populations

### v1

@@ 1 Introduction
Pulsars are immensely useful astrophysical probes for studying the underlying nuclear, plasma and gravitational physics that power a wide variety of observed phenomena (for recent reviews see, e.g., Kaspi & Kramer, 2016; Nättilä & Kajava, 2022). Originally discovered in the radio band (Hewish et al., 1968), they have since been shown to emit rotation-powered pulses across the electromagnetic spectrum (see Chapter 9 in Lorimer & Kramer, 2012, and references therein). Among the many high-energy instruments so far, Fermi Large Area Telescope (Fermi-LAT) has been revolutionary in opening the $\gamma-$ray window for pulsars (Atwood et al., 2009). Since its launch in 2008, Fermi-LAT has enabled the discovery of 294 pulsars that emit in $\gamma-$rays, with about a quarter of them so far only observable in $\gamma-$rays, making them 10% of all known pulsars Smith et al. (2023, hereafter 3PC). These sources are also the dominant Galactic component of $\gamma-$rays above 1 GeV. Furthermore, $\gamma-$ray luminosity of some pulsars has been observed to account for up to all of their total available spin-down energy (See Figure 24 in 3PC). This signifies the importance of studying $\gamma-$ray emission to better understand the high energy physics of pulsars. Understanding their Galactic population, such as their spatial distribution, total number and the number detectable by our telescopes, is important not only for stellar evolution studies but also to help us better constrain pulsar emission mechanisms. Population synthesis can be a viable tool for this purpose and will form the basis of the present study.
Pulsars broadly fall into two categories: millisecond pulsars (MSPs), with spin periods $P\lesssim 30$ ms, and canonical pulsars (CPs), with $P\gtrsim 30$ ms (for a recent review, see, e.g., Bhattacharyya & Roy, 2022). While accounting for the differences in period derivative, $\dot{P}$, between these two distinct populations, this simple cut-off in period separates the population into the older, recycled MSPs and the young CPs. Radio pulsar population syntheses have been used to constrain birth properties such as the birth rate and the period distribution for the CP and MSP pulsar population in multiple studies over the past half century (see, e.g., Gunn & Ostriker, 1970; Lyne et al., 1985; Narayan & Ostriker, 1990; Cordes & Chernoff, 1997; Faucher-Giguère & Kaspi, 2006; Gullón et al., 2014; Lorimer et al., 2015; Cieślar et al., 2020). Among other things, these studies have highlighted the difficulties in being able to constrain these distributions due to a range of different models being able to explain the observed population. Multi-wavelength population syntheses are one avenue to help remove these degeneracies, while also probing emission models at other wavelengths.
@@
More recently, Johnston et al. (2020) studied the highest energy and youngest CPs by imposing a cutoff on spin-down energy of $\dot{E}>10^{35}$ erg s-1. They assumed an empirical $\gamma-$ray luminosity law, as seen among Fermi-LAT CPs, of $L_{\gamma}\propto\sqrt{\dot{E}}$, while using the outer-gap (OG) emission geometry (Cheng et al., 1986). All of these works assumed a spin-down model of a vacuum retarded-dipole as first introduced by Deutsch (1955). Dirson et al. (2022) investigated all CPs and also used a better spin-down model based on the work of Contopoulos et al. (2014), which accounts for charged particles surrounding the pulsar magnetosphere in a force-free regime. They were able to test the empirical two dimensional (2D) plane luminosity relation for $\gamma-$ray luminosity proposed by Kalapotharakos et al. (2019), which includes a weak power-law dependence on the magnetic field, along with the $\sqrt{\dot{E}}$ term. The geometry used was based on the striped wind model (Pétri, 2011; Pétri & Mitra, 2021).
@@
Motivated by the goal of further understanding the nature of the unassociated Fermi-LAT sources, we approach the problem using population syntheses that are informed by our knowledge of the radio CP and MSP populations. We start by synthesizing a population consistent with the observed radio population of CPs and MSPs by fully accounting for selection effects from major blind radio surveys using PsrPopPy (Bates et al., 2014). We then test the effectiveness of multiple different $\gamma-$ray luminosity models using the sample of known $\gamma-$ray pulsars. Our aim is to investigate the efficacy of various $\gamma-$ray luminosity laws from the works of Bartels et al. (2018), Gonthier et al. (2018) and Kalapotharakos et al. (2023). By constraining the best $\gamma-$ray luminosity model, we also place limits on the number of $\gamma-$ray pulsars waiting to be discovered among the unassociated Fermi-LAT $\gamma-$ray sources. We finally assess the limitations of this work and suggest possible avenues for improvement in the future.
@@ 2 Fermi Catalog and $\gamma-$ray Pulsars
The latest Fermi catalog at the time of this work, 4FGL-DR4, contains 7194 sources (Abdollahi et al., 2022; Ballet et al., 2023). While the majority of these are blazars, the dominant source from our Galaxy is pulsars. Out of the 294 known $\gamma-$ray pulsars, 150 are CPs and 144 are MSPs. At the time of writing, 70 CPs are radio-quiet [^2] As defined in 3PC, they have radio flux density at 1400 MHz of at most 30 $\mu$Jy. whereas only 6 MSPs are radio-quiet. We exclude two of the radio-loud MSPs that are in globular clusters (J1824-2452A & J1823-3021A) and the double pulsar J0737-3039A from our analysis. We note that, while the double neutron star binary PSR J0737–3039A is also (by our $P<30$ ms definition) deemed a $\gamma-$ray MSP (Guillemot et al., 2013), its different evolutionary history means that we also excluded it from our sample of 141 $\gamma-$ray MSPs. For these 141 MSPs, 7 lack a reported $\dot{P}$ value in the catalog and are also excluded from our sample when comparing $P-\dot{P}$ plots or $\dot{P}$ distribution. These $\dot{P}$ lacking MSPs do contribute when comparing sizes of the samples. For CPs, we use the complete sample of 150. There are also 2428 sources in the catalog that remain unassociated with any astrophysical object and hence unidentified as an astronomical source. Out of these unidentified sources, 1061 are most likely to be pulsars due to the constraints on their observed properties. These constraints require the sources to have a log-parabolic spectrum, a spectral cut-off energy of 4 GeV and their fluxes do not significantly vary. Variability index (see section 3.6 of Nolan et al., 2012) is a proxy for the $\gamma-$ray flux variability from the mean value, with a value of greater than 24.7 indicating that there is a less than 1% chance of it being a steady source (for details, see Abdollahi et al., 2022). Trends seen in the observed $\gamma-$ray pulsar population guide these constraints. (Smith et al., 2023). A population synthesis of $\gamma-$ray pulsars should not predict the number of Fermi-LAT detectable $\gamma-$ray pulsars to be greater than the number of pulsar-like unidentified sources (henceforth referred to as PLUIDs) plus the number of known $\gamma-$ray pulsars.
@@ 3 Synthesizing Pulsars
Using a modified version of the pulsar population synthesis code PsrPopPy (Bates et al., 2014), we perform Monte Carlo simulations of pulsars observable by a set of radio surveys detailed below. We start by obtaining a birth period, age, magnetic field and magnetic inclination angle from appropriate distributions. Each pulsar is then placed in the model galaxy by first finding its radial position from the model galaxy’s center and then using this to place it in one of the spiral arms, following the procedure outlined in Appendix B of Faucher-Giguère & Kaspi (2006) and using updated parameters for the spiral arms as given in Yao et al. (2017). The pulsar’s $z$-component of spatial position is obtained through sampling a double-sided exponential distribution with a particular scale height. Next, each model pulsar is evolved in a model of the Galactic potential and its present day position in the Galactic plane is obtained. We used the adaption of the Carlberg & Innanen (1987) model of the Galactic potential given by Kuijken & Gilmore (1989).
@@ 3.2 Modeling considerations for millisecond pulsars
Most of the birth parameters for MSPs that we used are from the work of Lorimer et al. (2015, hereafter LEM+15). We made modifications to the log-normal magnetic field parameters through qualitative investigation to get the $P-\dot{P}$ plots that best match the observed sample. With a larger number of MSPs available now compared to LEM+15’s study, we decided to use this updated sample to derive a birth period distribution. We define the birth of an MSP as the time when the accretion from its companion stops and it becomes radio loud. Following Kalapotharakos et al. (2014) and GHF+18, in the ideal force-free magnetosphere, the period as a function of time
@@
Numerous $\gamma-$ray luminosity models have been proposed for MSPs using diverse methodologies by multiple researchers. We consider all of those which have been proposed in the past decade. First we have the purely statistical models from Bartels et al. (2018) which are derived from Bayesian model fitting on the observed sample. Their preferred model is a broken power law (BPL) but the log-normal (LN) distribution also had strong evidence so we test both of those. GHF+18 used the emission geometries from three different vacuum magnetosphere models of MSPs to find exponents for a power law $\gamma-$ray luminosity by doing population synthesis of radio and $\gamma-$ray MSPs. We tested all three of these models and their details are given in Table 3. Finally, we used the 2D and 3D FP of Kalapotharakos et al. (2019, 2022, 2023) since they are also applicable to MSPs. Since the 3D-plane relation requires that the $\gamma-$ray cutoff energy be known, we tested two different methods of reliably sampling this. We first tested the quadratic relationship, dependent on $\dot{E}$, first shown in Kalapotharakos et al. (2017), with updated parameters by fitting to the current sample of $\gamma-$ray MSPs. We also fit a Gaussian to the histogram of known cut-off energies. This model statistically agreed better with the observed sample than the quadratic model which is why we used it to obtain cutoff energies for our synthesized pulsars. For the three power-law models from GHF+18 and the fundamental planes, we also applied a noise parameter to account for the statistical uncertainty in measuring fluxes and distances used to estimate the luminosity of the observed sample. This noise component was drawn from a Gaussian with a mean of zero and standard deviation of 0.2 and added to the luminosity in log-space. We ran 10 iterations of this synthesis and then aggregated the results to obtain a mean number of detectable $\gamma-$ray MSPs. This number is presented, along with its standard deviation, in the plots comparing the expected number of detectable MSPs with the number suggested by each model. We note that our modeling of $\gamma-$ray MSPs does not invoke any beaming models, i.e., we explicitly assume all $\gamma-$ray MSPs are beaming towards Earth. We discuss the implications of this approach below.
@@ 3.3 Modeling considerations for canonical pulsars
where the time constant, $\tau_{D}=$ 5$\times$107 yr. As shown in JK17, this model is sufficient to produce a $P-\dot{P}$ plot which best matches the observed one, without the need for a decaying magnetic field. We verified their results using the birth assumptions they had made but chose to use the above mentioned parameters which are more reasonable and better substantiated. This still results in excellent overall matches in the $P-\dot{P}$ between the CP models and the observed population. We used the beaming model of Tauris & Manchester (1998).
After obtaining a model galaxy seeded with these CPs beaming towards us in radio, we calculate their $\gamma-$ray luminosities using the 2D FP, 3D FP and a heuristic model where $L_{\gamma}\propto\sqrt{\dot{E}}$. For the heuristic model we used the proportionality constant of $10^{16.5}~{}$erg s-1, as given in 3PC. These models also had the same noise component mentioned in the previous section added to them. We do not use a $\gamma-$ray beaming model for the 2D and 3D FP relations. For the heuristic model we initially employed a $\gamma-$ray beaming model based on the work of Song et al. (2024). This beaming model makes use of the $\dot{E}$ dependent beaming fractions originally highlighted by Johnston et al. (2020). However, as discussed further below, we modified it in a manner to get the best possible $P-\dot{P}$ distributions.
@@
Table 3: All the $\gamma-$ray luminosity models, their functional form and best fit parameters that were used in this work. The units for luminosity for all models are erg s-1 except for the GHF+18 models; OG, PSPC & TPC, for which they are eV s-1. For the FP models, B should be in units of Gauss, $\dot{E}$ should be erg s-1 and $\epsilon_{cut}$ should be MeV. In the GHF+18 models, $P$ is in units of ms and $\dot{P}$ is in units of $10^{-21}$s/s. $\dot{E}$ is in units of erg s-1 for the heuristic model. The range for luminosity in BPL and LN is [$10^{30},10^{37}$]. The FP, GHF+18 and heuristic models have a noise component drawn from a zero-mean Gaussian distribution with standard deviation 0.2 added to them in log space.

### v2

@@ 1 Introduction
Pulsars are immensely useful astrophysical probes for studying the underlying nuclear, plasma and gravitational physics that power a wide variety of observed phenomena (for recent reviews see, e.g., Kaspi & Kramer, 2016; Nättilä & Kajava, 2022). Originally discovered in the radio band (Hewish et al., 1968), they have since been shown to emit rotation-powered pulses across the electromagnetic spectrum (see Chapter 9 in Lorimer & Kramer, 2012, and references therein). Among the many high-energy instruments so far, Fermi Large Area Telescope (Fermi-LAT) has been revolutionary in opening the $\gamma-$ray window for pulsars (Atwood et al., 2009). Since its launch in 2008, Fermi-LAT has enabled the discovery of 294 pulsars that emit in $\gamma-$rays, making them 10% of all known pulsars Smith et al. (2023, hereafter 3PC). About a quarter of these Fermi-LAT discovered pulsars are so far only observable in $\gamma-$rays. These sources are also the dominant Galactic component of $\gamma-$rays above 1 GeV. Furthermore, $\gamma-$ray luminosity of some pulsars has been observed to account for up to all of their total available spin-down energy (See Figure 24 in 3PC). This signifies the importance of studying $\gamma-$ray emission to better understand the high energy physics of pulsars. Understanding their Galactic population, such as their spatial distribution, total number and the number detectable by our telescopes, is important not only for stellar evolution studies but also to help us better constrain pulsar emission mechanisms. Population synthesis can be a viable tool for this purpose and will form the basis of the present study.
Pulsars broadly fall into two categories: millisecond pulsars (MSPs), with spin periods $P\lesssim 30$ ms, and canonical pulsars (CPs), with $P\gtrsim 30$ ms (for a recent review, see, e.g., Bhattacharyya & Roy, 2022). While accounting for the differences in period derivative, $\dot{P}$, between these two distinct populations, this simple cut-off in period separates the population into the older, recycled MSPs and the young CPs. Radio pulsar population syntheses have been used to constrain properties, such as the birth rate and the period distribution, for the Milky Way’s underlying CP and MSP population in multiple studies over the past half century (see, e.g., Gunn & Ostriker, 1970; Lyne et al., 1985; Narayan & Ostriker, 1990; Cordes & Chernoff, 1997; Faucher-Giguère & Kaspi, 2006; Gullón et al., 2014; Lorimer et al., 2015; Cieślar et al., 2020). Among other things, these studies have highlighted the difficulties in being able to constrain these distributions due to a range of different models being able to explain the observed population. Multi-wavelength population syntheses are one avenue to help remove these degeneracies, while also probing emission models at other wavelengths.
@@
More recently, Johnston et al. (2020) studied the highest energy and youngest CPs by imposing a cutoff on spin-down energy of $\dot{E}>10^{35}$ erg s-1. They assumed an empirical $\gamma-$ray luminosity law, as seen among Fermi-LAT CPs, of $L_{\gamma}\propto\sqrt{\dot{E}}$, while using the outer-gap (OG) emission geometry (Cheng et al., 1986). All of these works assumed a spin-down model of a vacuum retarded-dipole as first introduced by Deutsch (1955). Dirson et al. (2022) investigated all CPs and also used a better spin-down model based on the work of Contopoulos et al. (2014), which accounts for charged particles surrounding the pulsar magnetosphere in a force-free regime. They were able to test the empirical two dimensional (2D) plane luminosity relation for $\gamma-$ray luminosity proposed by Kalapotharakos et al. (2019), which includes a weak power-law dependence on the magnetic field, along with the $\sqrt{\dot{E}}$ term. The geometry used was based on the striped wind model (Pétri, 2011; Pétri & Mitra, 2021). 2024A&A...691A.349S built upon the work of Dirson et al. (2022) by including a Galactic potential model in which the pulsars evolve over time and a minimum dispersion measure in their $\gamma-$ray detection model while still using the 2D plane luminosity relation and a binary Fermi-LAT sensitivity cutoff based on Galactic latitude. Their results showed that the striped wind model is able to reproduce the distribution of peak-separation seen in the $\gamma-$ray light curves of the observed $\gamma$-ray CP sample.
@@
Motivated by the goal of further understanding the nature of the unassociated Fermi-LAT sources, we approach the problem using population syntheses that are informed by our knowledge of the radio CP and MSP populations. We start by synthesizing a population consistent with the observed radio population of CPs and MSPs by accounting for selection effects from major blind radio surveys using PsrPopPy (Bates et al., 2014). We then test the effectiveness of multiple different $\gamma-$ray luminosity models using the sample of known $\gamma-$ray pulsars. Our aim is to investigate the efficacy of various $\gamma-$ray luminosity laws from the works of Bartels et al. (2018), Gonthier et al. (2018) and Kalapotharakos et al. (2023) in being able to produce a sample of pulsars compatible with the observed sample. Our work is not intended to constrain the best model. Instead, we elucidate the ways in which these models succeed and fail at reproducing the observed sample of pulsars after all the known multi-wavelength selection effects have been accounted for. We do select a fiducial $\gamma-$ray luminosity model to investigate more parameters and place limits on the number of $\gamma-$ray pulsars waiting to be discovered among the unassociated Fermi-LAT $\gamma-$ray sources. We finally assess the limitations of this work and suggest possible avenues for improvement in the future.
@@ 2 Fermi Catalog and $\gamma-$ray Pulsars
The latest Fermi catalog at the time of this work, 4FGL-DR4, contains 7194 sources (Abdollahi et al., 2022; Ballet et al., 2023). While the majority of these are blazars, the dominant source from our Galaxy is pulsars. Out of the 294 known $\gamma-$ray pulsars, 150 are CPs and 144 are MSPs. At the time of writing, of the observed $\gamma-$ray pulsars, 70 CPs are radio-quiet [^2] As defined in 3PC, they have radio flux density at 1400 MHz of at most 30 $\mu$Jy. whereas only 6 MSPs are radio-quiet. We exclude two of the radio-loud MSPs that are in globular clusters (J1824-2452A & J1823-3021A) from our analysis. We note that, while the double neutron star binary PSR J0737–3039A is also (by our $P<30$ ms definition) deemed a $\gamma-$ray MSP (Guillemot et al., 2013), its different evolutionary history means that we also excluded it from our sample of 141 $\gamma-$ray MSPs. For these 141 MSPs, 7 lack a reported $\dot{P}$ value in the catalog and are also excluded from our sample when comparing $P-\dot{P}$ plots or $\dot{P}$ distribution. These $\dot{P}$ lacking MSPs do contribute when comparing sizes of the samples. For CPs, we use the complete sample of 150. There are also 2428 sources in the catalog that remain unassociated with any astrophysical object and hence unidentified as an astronomical source. Out of these unidentified sources, 1061 are most likely to be pulsars due to the constraints on their observed properties. These constraints require the sources to have a log-parabolic spectrum, a spectral cut-off energy of 4 GeV and their fluxes do not significantly vary. Variability index (see section 3.6 of Nolan et al., 2012) is a proxy for the $\gamma-$ray flux variability from the mean value, with a value of greater than 24.7 indicating that there is a less than 1% chance of it being a steady source (for details, see Abdollahi et al., 2022). Trends seen in the observed $\gamma-$ray pulsar population guide these constraints. (Smith et al., 2023). A population synthesis of $\gamma-$ray pulsars should not predict the number of Fermi-LAT detectable $\gamma-$ray pulsars to be greater than the number of pulsar-like unidentified sources (henceforth referred to as PLUIDs) plus the number of known $\gamma-$ray pulsars.
@@ 3 Synthesizing Pulsars
Using a modified version of the pulsar population synthesis code PsrPopPy [^3] https://github.com/shawaizt/PsrPopPy-G (Bates et al., 2014), we perform Monte Carlo simulations of pulsars observable by a set of radio surveys listed in Table 1. We start by obtaining a birth period, age, magnetic field and magnetic inclination angle from appropriate distributions, as listed in Table 2. Each pulsar is then placed in the model galaxy by first finding its radial position from the model galaxy’s center and then using this to place it in one of the spiral arms, following the procedure outlined in Appendix B of Faucher-Giguère & Kaspi (2006) and using updated parameters for the spiral arms as given in Yao et al. (2017). The pulsar’s $z$-component of spatial position is obtained through sampling a double-sided exponential distribution with a particular scale height. Next, each model pulsar is evolved in a model of the Galactic potential and its present day position in the Galactic plane is obtained. We used the adaption of the Carlberg & Innanen (1987) model of the Galactic potential given by Kuijken & Gilmore (1989).
@@ 3.2 Modeling considerations for millisecond pulsars
Most of the birth parameters for MSPs that we used are from the work of Lorimer et al. (2015, hereafter LEM+15). With a larger number of MSPs available now compared to LEM+15’s study, we decided to use this updated sample to derive a birth period distribution. We define the birth of an MSP as the time when the accretion from its companion stops and it becomes radio loud. Following Kalapotharakos et al. (2014) and GHF+18, in the ideal force-free magnetosphere, the period as a function of time
@@
Numerous $\gamma-$ray luminosity models have been proposed for MSPs using diverse methodologies by multiple researchers. We consider all of those which have been proposed in the past decade. First we have the purely statistical models from Bartels et al. (2018) which are derived from Bayesian model fitting on the observed sample. Their preferred model is a broken power law (BPL) but the log-normal (LN) distribution also had strong evidence so we test both of those. GHF+18 used the emission geometries from three different vacuum magnetosphere models of MSPs to find exponents for a power law $\gamma-$ray luminosity by doing population synthesis of radio and $\gamma-$ray MSPs. We tested all three of these models and their details are given in Table 3. Finally, we used the 2D and 3D FP of Kalapotharakos et al. (2019, 2022, 2023) since they are also applicable to MSPs. Since the 3D-plane relation requires that the $\gamma-$ray cutoff energy is known, we tested two different methods of reliably sampling this. We first tested the quadratic relationship, dependent on $\dot{E}$, first shown in Kalapotharakos et al. (2017), with updated parameters by fitting to the current sample of $\gamma-$ray MSPs. We also fit a Gaussian to the histogram of known cut-off energies, $\epsilon_{\rm cut}$. Doing this in log-space, where $\epsilon_{\rm cut}$ is in units of Mev, we draw $\log_{10}\epsilon_{\rm cut}$ from a normal distribution with a mean of 3.458 and a standard deviation of 0.181. This model statistically agreed better with the observed sample than the quadratic model which is why we used it to obtain cutoff energies for our synthesized pulsars. For the three power-law models from GHF+18 and the fundamental planes, we also applied a noise parameter to account for the statistical uncertainty in measuring fluxes and distances used to estimate the luminosity of the observed sample. This noise component was drawn from a Gaussian with a mean of zero and standard deviation of 0.2 and added to the luminosity in log-space. We ran 10 iterations of this synthesis and then aggregated the results to obtain a mean number of detectable $\gamma-$ray MSPs. This number is presented, along with its standard deviation, in the plots comparing the expected number of detectable MSPs with the number suggested by each model. We note that our modeling of $\gamma-$ray MSPs does not invoke any beaming models, i.e., we explicitly assume all $\gamma-$ray MSPs are beaming towards Earth. We discuss the implications of this approach below.
@@ 3.3 Modeling considerations for canonical pulsars
where the time constant, $\tau_{D}=$ 5$\times$107 yr. As shown in JK17, this model is sufficient to produce a $P-\dot{P}$ plot which matches the observed one, without the need for a decaying magnetic field. We verified their results by looking at the $P-\dot{P}$ distribution obtained using their simplistic assumption that all pulsars are born with a period of 20 ms and the same $\dot{P}$. However, we chose to use birth assumptions which are more reasonable, such as a distribution of initial periods and a spin-down model. This still results in an excellent overall match to the $P-\dot{P}$ distribution between the CP models and the observed population. We used the beaming model of Tauris & Manchester (1998).
After obtaining a model galaxy seeded with these CPs beaming towards us in radio, we calculate their $\gamma-$ray luminosities using the 2D FP, 3D FP and a heuristic model where $L_{\gamma}\propto\sqrt{\dot{E}}$. For the heuristic model we used the proportionality constant of $10^{16.5}~{}$erg s-1, as given in 3PC. These models also had the same noise component mentioned in the previous section added to them. We do not use a $\gamma-$ray beaming model for the 2D and 3D FP relations. For the heuristic model we initially employed a $\gamma-$ray beaming model based on the work of Song et al. (2024). This beaming model makes use of the $\dot{E}$ dependent beaming fractions originally highlighted by Johnston et al. (2020). However, as discussed further below, we modified it to get a $P-\dot{P}$ distribution which is qualitatively most similar to the observed distribution.
@@
Table 3: All the $\gamma-$ray luminosity models, their functional form and best fit parameters that were used in this work. The units for luminosity for all models are erg s-1 except for the GHF+18 models; OG, PSPC & TPC, for which they are eV s-1. For the FP models, B should be in units of Gauss, $\dot{E}$ should be erg s-1 and $\epsilon_{\rm cut}$ should be MeV. In the GHF+18 models, $P$ is in units of ms and $\dot{P}$ is in units of $10^{-21}$s/s. $\dot{E}$ is in units of erg s-1 for the heuristic model. The range for luminosity in BPL and LN is [$10^{30},10^{37}$]. The FP, GHF+18 and heuristic models have a noise component drawn from a zero-mean Gaussian distribution with standard deviation 0.2 added to them in log space.